Exercise 2-15 Plantwide and Departmental Predetermined Overhead Rates; Job Costs [LO2-1, LO2-2, LO2-3, LO2-4]

[The following information applies to the questions displayed below.]

Delph Company uses a job-order costing system and has two manufacturing departments—Molding and Fabrication. The company provided the following estimates at the beginning of the year:

  

Molding Fabrication Total
Machine-hours 33,000 43,000 76,000
Fixed manufacturing overhead costs $ 800,000 $ 200,000 $ 1,000,000
Variable manufacturing overhead cost per machine-hour $ 5.70 $ 5.70

  

During the year, the company had no beginning or ending inventories and it started, completed, and sold only two jobs—Job D-70 and Job C-200. It provided the following information related to those two jobs:

  

Job D-70: Molding Fabrication Total
Direct materials cost $ 379,000 $ 330,000 $ 709,000
Direct labor cost $ 220,000 $ 130,000 $ 350,000
Machine-hours 24,000 9,000 33,000

  

Job C-200: Molding Fabrication Total
Direct materials cost $ 210,000 $ 260,000 $ 470,000
Direct labor cost $ 170,000 $ 240,000 $ 410,000
Machine-hours 9,000 34,000 43,000

Delph had no underapplied or overapplied manufacturing overhead during the year.

Exercise 2-15 Part 1

Required:

1. Assume Delph uses a plantwide predetermined overhead rate based on machine-hours.

a. Compute the plantwide predetermined overhead rate.

b. Compute the total manufacturing cost assigned to Job D-70 and Job C-200.

c. If Delph establishes bid prices that are 140% of total manufacturing cost, what bid prices would it have established for Job D-70 and Job C-200?

d. What is Delph’s cost of goods sold for the year?

Brewer 8e Rechecks 2018-08-21
Answer #1

1a) Plantwide overhead rate = 1000000/76000+5.70 = 18.86 per MH

1b) Total manufacturing cost

Job D-70 Job C-200
Direct material 709000 470000
Direct labor 350000 410000
Overhead 622380 810980
Total manufacturing cost 1681380 1690980

1c) Bid price

Job D-70 = 1681380*1.4 = 2353932

Job C-200 = 1690980*1.4 = 2367372

1d) Cost of goods sold = 1681380+1690980 = 3372360